4. A researcher claims that the distribution of favorite pizza toppings among teenagers is as shown below. Topping Cheese Pepperoni Sausage Mushrooms Onions Frequency | 41% 25% 15% 10% 9%
200 teenagers are randomly selected and asked what their favorite pizza topping is. The results are shown below. Topping Cheese Peperoni Sausage Mushrooms Onions 30 52 Results (n 20078 (a) Set up the appropriate Ho and Ha (b) Calculate the test statistic. (c) Find the P-value and make your conclusion at α 25 15 .05.
Answer #1

Solution:-

State the hypotheses. The first step is to state the null hypothesis and an alternative hypothesis.

Null hypothesis: The distribution of favorite pizza toppings among teenagers is as per given distribution.

Alternative hypothesis: At least one of the proportions in the null hypothesis is false.

Formulate an analysis plan. For this analysis, the significance level is 0.05. Using sample data, we will conduct a chi-square goodness of fit test of the null hypothesis.

Analyze sample data. Applying the chi-square goodness of fit test to sample data, we compute the degrees of freedom, the expected frequency counts, and the chi-square test statistic. Based on the chi-square statistic and the degrees of freedom, we determine the P-value.

DF = k - 1 = 5 - 1
D.F = 4
(Ei) = n * pi
Observed E(x Expected Num ((Or r,c)21 Er 82 0.19512195121 Peperoni Sausage Mushrooms Onion Sum 78 52 30 25 15 200 0.41 0.25 0
Er,c
X2 = 2.025

where DF is the degrees of freedom, k is the number of levels of the categorical variable, n is the number of observations in the sample, Ei is the expected frequency count for level i, Oi is the observed frequency count for level i, and X2 is the chi-square test statistic.

The P-value is the probability that a chi-square statistic having 4 degrees of freedom is more extreme than 2.025.

We use the Chi-Square Distribution Calculator to find P(X2 > 2.025) = 0.7311.

Interpret results. Since the P-value (0.7371) is greater than the significance level (0.05), we have to accept the null hypothesis.

From the above test we can conclude that the distribution of favorite pizza toppings among teenagers is as per given distribution.